A photon emitted by an excited atom has an energy of 2.5 eV. What is its frequency?
Answer #1

Solution)

Given,

Energy E=2.5*(1.6*10^-19)V= 4*10^-19 V

We know,

E=h*f

Substitute values,

Frequency, f=(4*10^-19)/(6.6*10^-34)= 6.06*10^14 Hz (Ans)

========